b where v7 = (2, -1), wł = (-1,2), b7 = 2. Find real numbers c and d so that cv + dw (1,0).
Answer #1

Solution : T = (2-1) WT = (-1, 2) , 67 = (1,0) cvt dw = b (2)+ (%) = (0) 2c + ad 2c - d -ct ad ac - d = 1 -ct ad=0 2 Solve an